OpenStudy (anonymous):

Factor the trinomial: 3x2 + 11x + 6

OpenStudy (anonymous):

i suck at math too but do you want my help anyway?

OpenStudy (anonymous):

haha yes

OpenStudy (anonymous):

these guys will prolly get the answer before i get so i'll wait....:)

OpenStudy (anonymous):

\[3x^2 + 11x + 6 = 3x^2 + 9x + 2x + 6\] Can you now guess what will be the factor which will be taken out ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

me niether...

OpenStudy (anonymous):

@dpaInc doing great job go ahead with your work...

OpenStudy (anonymous):

-2/3 and -3

OpenStudy (anonymous):

can someone have a phone and can text me?i cant do this im just a 14 year old girl!!lol

OpenStudy (anonymous):

Don't give out your number kid!

OpenStudy (anonymous):

thats not one of the choices for my online summerschool thing

OpenStudy (anonymous):

wht r the choices?

OpenStudy (anonymous):

A. (3x + 2)(x + 3) B. (3x + 3)(x + 2) C. (3x + 1)(x + 6) D. (3x + 6)(x + 1)

myininaya (myininaya):

Hey you guys! Please try to help the person get the answer on their own. Giving the answer doesn't help. Also it is totally not safe to give out private info. Thanks,

OpenStudy (anonymous):

^lets not listen to her we need answers

myininaya (myininaya):

Oh no! You will listen to me if you don't want to be suspended. It is also against the CoC to just look for answers.

OpenStudy (anonymous):

leave it at last 3rd step

myininaya (myininaya):

Lets try to teach this person please.

OpenStudy (anonymous):

Their actually right, dont give exactly the answer. We will not learn if we give the answer and Im sorry if the other question i give all the answer.. I didnt think about that today..

myininaya (myininaya):

3x^2+11x+6 ax^2+bx+c The trick I like to use is: Find two factors of a*c that have product a*c and sum b

myininaya (myininaya):

So we need to find two factors of 3*6 that have product 18 and sum 11

myininaya (myininaya):

Can you @isuckatmath419 think of two such numbers?

OpenStudy (anonymous):

i can think of 3 is that one?

myininaya (myininaya):

What two numbers can you multiply to get 18 and add to get 11 is my question

OpenStudy (anonymous):

@myininaya are you a teacher? cuz if you are, mybe if i hav a question for math i might ask u.. u sound like a teacher for me..XD (sorry about that)

myininaya (myininaya):

Yes I teach mathematics.

OpenStudy (anonymous):

i cant think of the two numbers

myininaya (myininaya):

Hint: What is 9(2) and 9+2?

OpenStudy (anonymous):

ohhhh haha its 9 and 2

myininaya (myininaya):

lol yes! :) Now once you find those numbers you replace your middle term 11x with 9x+2x And this is still the same expression just written a little different 3x^2+9x+2x+6

myininaya (myininaya):

Then you factor by grouping

myininaya (myininaya):

What factors do 3x^2 and 9x have in common?

OpenStudy (anonymous):

@myininaya Oooh Ok, thank u. At least I know one teacher now.. or something.. well i know one too who is good but his/her a student... cuz i can tell that Im not really good at Math..

OpenStudy (anonymous):

have a 3 in common

myininaya (myininaya):

And what factor by grouping means is we will look at the first two terms together and the last two terms together like this: (3x^2+9x)+(2x+6) And then we will "pull out" what they have in common from each grouping And they also have the factor x in common too, right?

myininaya (myininaya):

So that means we can factor 3x from the first two terms like 3x(....+....)+(2x+6)

myininaya (myininaya):

But what goes in those two blanks once we "pull out" 3x

myininaya (myininaya):

Or in other words divide both 3x^2 and 9x by 3x since 3x is the thing you "pulled out"

OpenStudy (anonymous):

is that 1x?

myininaya (myininaya):

What is \[\frac{3x^2}{3x}=? \text{ and } \frac{9x}{3x}=?\]

OpenStudy (anonymous):

im not sure,i thought maybe 1x

myininaya (myininaya):

\[\frac{3 x x}{3 x} =? \text{ and } \frac{ 3 \cdot 3 x}{3 x}=?\]

myininaya (myininaya):

Is it obvious now?

myininaya (myininaya):

oh yes for that one it is 1x but what about the other

OpenStudy (anonymous):

is it 6x?

myininaya (myininaya):

And by the way 1x=x so we can just say x instead of 1x but 1x is fine but you won't see your answer with 1x you will see it with just the x well the 3x's cancel right leaving you just with ....

myininaya (myininaya):

what is 9/3?

OpenStudy (anonymous):

its the same as 6/3?

myininaya (myininaya):

nope 9/3 is 3 because 3*3=9

myininaya (myininaya):

6/3 is 2 because 2*3=6 but the question was what is 9/3 not 6/3

myininaya (myininaya):

So we have 3x(x+3)+(2x+6) You try the second pair of terms

myininaya (myininaya):

Factor 2x+6

myininaya (myininaya):

What factors does 2x and 6 have in common?

OpenStudy (anonymous):

oh 2?

myininaya (myininaya):

right! :) So if we factor out 2 from the second two terms we have 3x(x+3)+2(...+...) But what goes in those blanks

myininaya (myininaya):

Just divide 2x by 2 and just divide 6 by 2

OpenStudy (anonymous):

so is that 1 and 3?

myininaya (myininaya):

Close

myininaya (myininaya):

2x/2 leaves us with also that x right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

So we have 3x(x+3)+2(x+3) But omg right! We have two terms left and they both have what common factor?

myininaya (myininaya):

And when I say two terms left I mean the first term is 3x(x+3) and the second term is 2(x+3)

myininaya (myininaya):

Both of these terms have the common factor _____?_____

myininaya (myininaya):

You are almost there! You can do it @isuckatmath419 :)

OpenStudy (anonymous):

umm idk is it 1 3?

myininaya (myininaya):

Ok just so you know the first term has the following factors: 3, x , (x+3) The second term has the following factors: 2, (x+3) They both have what factor in common?

myininaya (myininaya):

They don't have 2 in common They don't have 3 in common They don't have x in common But they do have ...... in common

OpenStudy (anonymous):

is it they got + in common

myininaya (myininaya):

they have the factor (x+3) in common

myininaya (myininaya):

Do you see this reply: "Ok just so you know the first term has the following factors: 3, x , (x+3) The second term has the following factors: 2, (x+3) They both have what factor in common?" The first term had three factors, the 3 ,the x , the (x+3) The second term had two factors, the 2 , the (x+3) The (x+3) is a factor in both of these terms

myininaya (myininaya):

So anyways we are trying to factor this 3x(x+3)+2(x+3) And we said (x+3) is the factor that both terms have in common

myininaya (myininaya):

So we can "pull out" (x+3)

myininaya (myininaya):

(x+3)(___+___) Just divide your first term 3x(x+3) by (x+3) and Just divide your second term 2(x+3) by (x+3) to figure out what goes in those blanks and then you will have your answer

myininaya (myininaya):

What did you get when you divided 3x(x+3) by (x+3) ?

OpenStudy (anonymous):

im trying im not good at it

myininaya (myininaya):

What do you get when you divide 3xy by y?

OpenStudy (anonymous):

3xy*?

myininaya (myininaya):

Well the y's cancel You have factor y on top and factor y on bottom so they cancel

myininaya (myininaya):

3xy/y=3x So 3x(x+3)/(x+3) is also the same 3x You have the factor (x+3) on top and the factor (x+3) on bottom they cancel

myininaya (myininaya):

Now 3x goes in the first blank so so far we have 3x(x+3)+2(x+3) (x+3)(___+___) (x+3)(3x+___)

myininaya (myininaya):

now what is 2(x+3)/(x+3) =?

myininaya (myininaya):

You have the factor (x+3) on top and bottom so what are you gonna do what the (x+3)'s ?

myininaya (myininaya):

@isuckatmath419 can you do this division 2(x+3)/(x+3) =?